The floor of a building consists of 3000 tiles which are rhombus shaped and each of its diagonals are 45 cm and 30 cm in length. Find the total cost of polishing the floor, if the cost per m2is Rs. 4.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Length of one diagonal, d1= 45 cm and d2= 30 cm
∵Area of one tile = (1⁄2)d1d2
= (1⁄2)×45×30= 675
Area of one tile is 675 cm^2
Area of 3000 tiles is
= 675×3000 = 2025000 cm^2
= 2025000/10000
= 202.50m^2[∵1m^2= 10000 cm^2]
∵Cost of polishing the floor per sq. meter = 4
Cost of polishing the floor per 202.50 sq. meter = 4×202.50 = 810
Hence the total cost of polishing the floor is Rs. 810.

Thanks